# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v1

@@ 6 CONCLUSION
In this paper, we established existence of Nash equilibria in continuous games with playerwise concave utilities and shared, playerwise concave constraints. We proposed an independent log barrier regularized gradient method and proved its convergence to constrained Nash equilibria in potential games with shared coupling constraints.
Our work suggests several directions for future work: Can equilibrium existence and computation be characterized when utilities or constraints lack playerwise concavity, as in Markov games? Do our results extend to settings with stochastic zeroth- or first-order feedback? And finally, one may ask whether our approach can be applied to learning weaker notions of equilibria in general-sum games, or constrained Nash equilibria in other tractable subclasses of playerwise concave games, such as zero-sum or monotone games.
@@ Appendix A Overview of Notation
$\Phi(x)$ $\coloneqq\quad$ potential function for $m$-player game in Section 4.1
@@ Appendix D Proof of Existence Result
Suppose Assumption 1 holds, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. For any $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, the set
@@
Suppose Assumption 2 holds, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. Then, for all $i\in[m]$, the set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is continuous (i.e. both upper and lower semicontinuous) over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$.
@@
Let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$, which exists as $\mathcal{C}\not$$\emptyset$. The proof is divided into two parts:
  1. 1)
    Contractibility of $\mathcal{C}^{\prime}$: For a compact set $X\subset\mathbb{R}^{n}$, its centroid is defined as
@@
    Regarding 1)(a), observe that each $h_{i}$ is well-defined: By definition, we have $x_{i}\in\mathcal{C}^{\prime}_{i}(x_{-i})$. By convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$, as shown in Lemma 1, it holds that $\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}_{i}(x_{-i})$. Thus for all $t\in[0,1]$, again by convexity, $(1-t)x_{i}+t\,\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}$. As $h_{i}$ keeps $x_{-i}$ fixed, we have $h_{i}(t,x)\in\mathcal{C}^{\prime}$ for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$. Then, by definition, $H_{1}$ is well-defined, and the same follows for each $H_{i}$ inductively. We conclude that $H$ is well-defined, as it is piecewise equivalent to the $H_{i}$’s.
@@
    To show 1)(b), note that by Lemma 2, $\mathcal{C}^{\prime}_{i}(x_{-i})$ is continuous in $x_{-i}$. Then, continuity of each $h_{i}$ in $(t,x)$ follows together with the continuity of the centroid function $\overline{c}(\cdot)$. Similarly as in (a), the recursive definitions of each $H_{i}$ and the piecewise definition of $H$ lets us conclude that $H$ is indeed continuous in $(t,x)$. In particular, we point out that continuity at the boundary of the pieces is ensured, as for all $i\in[m]\setminus\left\{1\right\}$ and all $x\in\mathcal{C}^{\prime}$, by definition, $\lim_{t\to{(i-1)/m}^{+}}H_{i}(t,x)=H_{i-1}(\varphi^{-1}_{i-1}(1),x)$.
@@
    For 1)(c), by definition we have $H(0,x)=x$ for all $x\in\mathcal{C}$. To argue about $H(1,\cdot)$, let $\mathcal{C}^{\prime(0)}\coloneqq\mathcal{C}^{\prime}$ and, for $i\in[m]$, let $\mathcal{C}^{\prime(i)}\subseteq\mathcal{C}^{\prime}$ denote the image of the function $x\mapsto H_{i}(1,x)$. Thus, $\mathcal{C}^{\prime(i)}\coloneqq\{H_{i}(\varphi^{-1}_{i}(1),x)\mid x\in\mathcal{C}^{\prime}\}$ which due to the recursive definition of $H_{i}$, for all $i\in[m]$, is equal to $\{h_{i}(1,x)\mid x\in\mathcal{C}^{\prime(i-1)}\}$. Let $d_{i}\coloneqq\text{dim}(\mathcal{X}_{i})$ and notice that $\text{dim}(\mathcal{C}^{\prime(0)})\leq\sum_{i\in[m]}d_{i}$. Next, we have $\mathcal{C}^{\prime(1)}=\{(\overline{c}(\mathcal{C}_{1}^{\prime}(x_{-i})),x_{-i})\mid\exists x_{1}\in\mathcal{X}_{1}\text{ s.t.}\;(x_{1},x_{-1})\in\mathcal{C}^{\prime(0)}\}$, meaning $\mathcal{C}^{\prime(1)}$ is can be described as a parameterized set with parameters from a set of dimension at most $\sum_{i=2}^{m}d_{i}$. Hence $\text{dim}(\mathcal{C}^{\prime(1)})\leq\sum_{i=2}^{m}d_{i}$. Similarly, we can argue that for all $1<i\leq[m]$, $\text{dim}(\mathcal{C}^{\prime(i)})\leq\sum_{j=i+1}^{m}d_{i}$, since $\mathcal{C}^{\prime(i)}$ can be described as a parameterized set where coordinates corresponding to the first $i$ players are determined as a function of the remaining players $i+1$ to $m$. In particular, for the image of $x\mapsto H(1,x)$, namely the set $\mathcal{C}^{\prime(m)}$, we have therefore have $\text{dim}(\mathcal{C}^{\prime(m)})\leq 0$. Finally, in order to argue that $\mathcal{C}^{\prime(m)}$ is connected, we recall that the image of any continuous function with connected domain is connected. Since $\mathcal{C}^{\prime}$ was chosen as a connected component, by continuity of $x\mapsto H(1,x)$, $\mathcal{C}^{\prime(m)}$ is therefore connected. As a connected set of dimension 0, $\mathcal{C}^{\prime(m)}$ can only contain a single point $\hat{x}$, meaning $H(1,x)=\hat{x}$ for all $x\in\mathcal{C}$.
  2. 2)
@@
    Fixed point argument: For any $\epsilon>0$, define the perturbed game $\Gamma^{\epsilon}$ with regularized utilities $u_{i}^{\epsilon}(x)\coloneqq u_{i}(x)-\epsilon\lVert x_{i}\rVert^{2}$; constraints are left unchanged. Observe that $\Gamma^{\epsilon}$ has playerwise strongly concave utilities. We first show that for any $\epsilon>0$, $\Gamma^{\epsilon}$ has a constrained Nash equilibrium in $\mathcal{C}^{\prime}$. Then we obtain the existence result for the original game via a limit argument.
@@
    Since $\mathcal{C}^{\prime}$ is compact and contractible, by Fact 3.2, there exists $x^{\star}\in\mathcal{C}^{\prime}$ such that $x^{\star}=\Psi(x^{\star})$. As each $\Psi_{i}$ leaves the strategies of players other than $i$ unchanged, this implies that $\Psi_{i}(x^{\star})=x^{\star}$ for all $i\in[m]$. Suppose $x^{\star}$ is not a constrained Nash equilibrium. Then for some $i\in[m]$, there exists $x_{i}^{\prime}$ such that $\hat{x}\coloneqq(x_{i}^{\prime},x^{\star}_{-i})\in\mathcal{C}$ and $u_{i}(\hat{x})>u_{i}(x^{\star})$. We observe that by playerwise concavity of constraints, the line segment $x^{\star},\hat{x}]\subset\mathcal{C}$ satisfies all constraints, meaning $x^{\star}$ and $\hat{x}$ lie in the same connected component of $\mathcal{C}$; thus $\hat{x}\in\mathcal{C}^{\prime}\subseteq\mathcal{C}$. We have reached a contradiction with $x^{\star}_{i}$ being the maximizer in the definition of $\Psi_{i}(x^{\star})$, hence $x^{\star}$ is a constrained Nash equilibrium of $\Gamma^{\epsilon}$.
@@
    i.e., we have shown that $x^{\star}$ is a constrained Nash equilibrium of the original unperturbed game.
@@ E.2 Technical lemma
Appendix F Proofs for Section 4.1
In this section, we provide the full proofs for the results of Section 4.1.
@@ F.1 Mixed-strategy potential games
In Example 3 we claimed that when considering mixed strategies over a finite action potential game, the potential structure is lifted to the continuous game over compact convex set of mixed strategies. Here we state the corresponding proposition which is a direct consequence of Monderer and Shapley, (1996), Lemma 2.10.
@@
Let $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ with each $\mathcal{A}_{i}$ finite, and suppose $\widetilde{u}_{i}:\mathcal{A}\to[0,1]$ are player utilities such that there exists a potential function $\phi:\mathcal{A}\to\mathbb{R}$ satisfying
$\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})\quad\text{for all }i\in[m],\,a\in\mathcal{A},\,a_{i}^{\prime}\in\mathcal{A}_{i}.$
Then, the continuous game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$ and expected utilities
@@
is a potential game with potential function $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$.
@@ F.2 Technical lemmas
It is immediate from Assumption 6 that if we have a game with utilities $\{u^{\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime}$, and another game with utilities $\{u^{\prime\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime\prime}$, then $\Phi^{\prime}+\Phi^{\prime\prime}$ is a potential function for the game with utilities $\{u^{\prime}_{i}+u^{\prime\prime}_{i}\}_{i\in[m]}$. The statement hence follows by observing that the sum of log barrier terms forms an identical interest game and hence $\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for it. ∎
@@ F.4 Lower bounding the stepsize
Finally, we are ready to prove the main convergence result for the constrained potential game setting.
Theorem 3.
Let $\epsilon>0$ and set $\eta=\epsilon$. Let Assumptions 1, 3, 4, 6, and 5 hold with $\rho\geq\eta$. For stepsizes $\{\gamma^{(t)}\}_{t=0}^{T-1}$ chosen as in (6), suppose we perform $T=\mathcal{O}(\epsilon^{-3})$ rounds of simultaneous updates as in (1). Then, there exists $0\leq t\leq T$ such that $x^{(t)}$ is a constrained $\epsilon$-approximate Nash equilibrium.

### v2

@@ 6 CONCLUSION
In this paper, we established existence of Nash equilibria in continuous action games with playerwise concave utilities and shared, playerwise concave constraints. Furthermore, we proposed an independent log barrier regularized gradient method and proved its convergence to constrained Nash equilibria in potential games with shared coupling constraints.
Our findings suggest several directions for future work. Can equilibrium existence and computation be characterized when utilities or constraints lack playerwise concavity, as in Markov games? Does our algorithmic approach extend to settings with stochastic zeroth- or first-order feedback? And finally, we ask whether our approach can be applied to learning equilibria in other tractable subclasses of playerwise concave games, such as zero-sum or monotone games.
@@ Appendix A Overview of Notation
$\operatorname{supp}(\mathcal{C}_{i})$ $\coloneqq\quad\left\{x_{-i}\in\mathcal{X}_{-i}\mid\exists x_{i}\in\mathcal{X}_{i}\text{ s.t.\ }(x_{i},x_{-i})\in\mathcal{C}\right\}$
@@
$\Phi(x)$ $\coloneqq\quad$ potential function for $m$-player game in Section 4
@@ C.1 Contractibility
C.2 Continuity for set-valued functions
We next recall continuity concepts for set-valued functions, which will be used when discussing continuity properties of feasible response mappings.
Definition 3.
Let $Y,Z\subseteq\mathbb{R}^{n}$, and let $F:Y\to\mathcal{P}(Z)$ be a set-valued function.
  1. 1.
    The function $F$ is upper semicontinuous at a point $y_{0}\in Y$ if and only if for any open subset $V\subseteq Z$ with $F(y_{0})\subseteq V$ there exists a neighborhood $U(y_{0})$ of $y_{0}$, such that $F(y)\subseteq V$ for all $y\in U(y_{0})$.
  2. 2.
    The function $F$ is lower semicontinuous at a point $y_{0}\in Y$ if and only if for any open subset $V\subseteq Z$ with $F(y_{0})\cap V\not$$\emptyset$ there exists a neighborhood $U(y_{0})$ of $y_{0}$, such that $F(y)\cap V\not$$\emptyset$ for all $y\in U(y_{0})$.
  3. 3.
    The function $F$ is continuous if and only if it is both upper and lower semicontinuous.
To introduce a notion of Lipschitz continuity for set-valued functions, we first define the Hausdorff distance between two nonempty sets $Y,Z\subseteq\mathbb{R}^{n}$ as
$d_{H}(Y,Z)\coloneqq\max\left\{\sup_{y\in Y}\text{dist}(y,Z),\;\sup_{z\in Z}\text{dist}(z,Y)\right\}$
where $\text{dist}(y,Z)\coloneqq\inf_{z\in Z}\left\lVert y-z\right\rVert$.
Definition 4.
Let $Y,Z\subseteq\mathbb{R}^{n}$ and $L\geq 0$. The set-valued function $F:Y\to\mathcal{P}(Z)$ is said to be $L$-Lipschitz continuous with respect to the Hausdorff distance if for all $y_{1},y_{2}\in Y$, we have
$d_{H}(F(y_{1}),F(y_{2}))\leq L\left\lVert y_{1}-y_{2}\right\rVert.$
C.3 Berge’s maximum theorem
@@ Appendix D Proof of Existence Result
Suppose Assumption 1 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. For any $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, the set
@@
Suppose Assumption 2 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. Then, for all $i\in[m]$, the set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is continuous (i.e. both upper and lower semicontinuous) over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$.
@@
For a compact set $X\subset\mathbb{R}^{n}$, its centroid is defined as
@@
Regarding (a), observe that each $h_{i}$ is well-defined: By definition, we have $x_{i}\in\mathcal{C}^{\prime}_{i}(x_{-i})$. By convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$, as shown in Lemma 1, it holds that $\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}_{i}(x_{-i})$. Thus for all $t\in[0,1]$, again by convexity, $(1-t)x_{i}+t\,\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}$. As $h_{i}$ keeps $x_{-i}$ fixed, we have $h_{i}(t,x)\in\mathcal{C}^{\prime}$ for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$. Then, by definition, $H_{1}$ is well-defined, and the same follows for each $H_{i}$ inductively. We conclude that $H$ is well-defined, as it is piecewise equivalent to the $H_{i}$’s.
@@
To show (b), note that by Lemma 2, $\mathcal{C}^{\prime}_{i}(x_{-i})$ is continuous in $x_{-i}$. Then, continuity of each $h_{i}$ in $(t,x)$ follows together with the continuity of the centroid function $\overline{c}(\cdot)$. Similarly as in (a), the recursive definitions of each $H_{i}$ and the piecewise definition of $H$ lets us conclude that $H$ is indeed continuous in $(t,x)$. In particular, we point out that continuity at the boundary of the pieces is ensured, as for all $i\in[m]\setminus\left\{1\right\}$ and all $x\in\mathcal{C}^{\prime}$, by definition, $\lim_{t\to{(i-1)/m}^{+}}H_{i}(t,x)=H_{i-1}(\varphi^{-1}_{i-1}(1),x)$.
@@
For (c), by definition we have $H(0,x)=x$ for all $x\in\mathcal{C}$. To argue about $H(1,\cdot)$, let $\mathcal{C}^{\prime(0)}\coloneqq\mathcal{C}^{\prime}$ and, for $i\in[m]$, let $\mathcal{C}^{\prime(i)}\subseteq\mathcal{C}^{\prime}$ denote the image of the function $x\mapsto H_{i}(1,x)$. Thus, $\mathcal{C}^{\prime(i)}\coloneqq\{H_{i}(\varphi^{-1}_{i}(1),x)\mid x\in\mathcal{C}^{\prime}\}$ which due to the recursive definition of $H_{i}$, for all $i\in[m]$, is equal to $\{h_{i}(1,x)\mid x\in\mathcal{C}^{\prime(i-1)}\}$. Let $d_{i}\coloneqq\text{dim}(\mathcal{X}_{i})$ and notice that $\text{dim}(\mathcal{C}^{\prime(0)})\leq\sum_{i\in[m]}d_{i}$. Next, we have $\mathcal{C}^{\prime(1)}=\{(\overline{c}(\mathcal{C}_{1}^{\prime}(x_{-i})),x_{-i})\mid\exists x_{1}\in\mathcal{X}_{1}\text{ s.t.}\;(x_{1},x_{-1})\in\mathcal{C}^{\prime(0)}\}$, meaning $\mathcal{C}^{\prime(1)}$ is can be described as a parameterized set with parameters from a set of dimension at most $\sum_{i=2}^{m}d_{i}$. Hence $\text{dim}(\mathcal{C}^{\prime(1)})\leq\sum_{i=2}^{m}d_{i}$. Similarly, we can argue that for all $1<i\leq[m]$, $\text{dim}(\mathcal{C}^{\prime(i)})\leq\sum_{j=i+1}^{m}d_{i}$, since $\mathcal{C}^{\prime(i)}$ can be described as a parameterized set where coordinates corresponding to the first $i$ players are determined as a function of the remaining players $i+1$ to $m$. In particular, for the image of $x\mapsto H(1,x)$, namely the set $\mathcal{C}^{\prime(m)}$, we have therefore have $\text{dim}(\mathcal{C}^{\prime(m)})\leq 0$. Finally, in order to argue that $\mathcal{C}^{\prime(m)}$ is connected, we recall that the image of any continuous function with connected domain is connected. Since $\mathcal{C}^{\prime}$ was chosen as a connected component, by continuity of $x\mapsto H(1,x)$, $\mathcal{C}^{\prime(m)}$ is therefore connected. As a connected set of dimension 0, $\mathcal{C}^{\prime(m)}$ can only contain a single point $\hat{x}$, meaning $H(1,x)=\hat{x}$ for all $x\in\mathcal{C}$. ∎
[…]
Proof.
Let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be any connected component of $\mathcal{C}$, which exists as $\mathcal{C}\not$$\emptyset$.
@@
For any $\epsilon>0$, define the perturbed game $\Gamma^{\epsilon}$ with regularized utilities $u_{i}^{\epsilon}(x)\coloneqq u_{i}(x)-\epsilon\lVert x_{i}\rVert^{2}$; constraints are left unchanged. Observe that $\Gamma^{\epsilon}$ has playerwise strongly concave utilities. We first show that for any $\epsilon>0$, $\Gamma^{\epsilon}$ has a constrained Nash equilibrium in $\mathcal{C}^{\prime}$. Then we obtain the existence result for the original game via a limit argument.
@@
Since $\mathcal{C}^{\prime}$ is compact and contractible (by Lemma 3.2), it follows from Fact 3.2 that there exists $x^{\star}\in\mathcal{C}^{\prime}$ such that $x^{\star}=\Psi(x^{\star})$. As each $\Psi_{i}$ leaves the strategies of players other than $i$ unchanged, this implies that $\Psi_{i}(x^{\star})=x^{\star}$ for all $i\in[m]$. Suppose $x^{\star}$ is not a constrained Nash equilibrium. Then for some $i\in[m]$, there exists $x_{i}^{\prime}$ such that $\hat{x}\coloneqq(x_{i}^{\prime},x^{\star}_{-i})\in\mathcal{C}$ and $u_{i}(\hat{x})>u_{i}(x^{\star})$. We observe that by playerwise concavity of constraints, the line segment $x^{\star},\hat{x}]\subset\mathcal{C}$ satisfies all constraints, meaning $x^{\star}$ and $\hat{x}$ lie in the same connected component of $\mathcal{C}$; thus $\hat{x}\in\mathcal{C}^{\prime}\subseteq\mathcal{C}$. We have reached a contradiction with $x^{\star}_{i}$ being the maximizer in the definition of $\Psi_{i}(x^{\star})$, hence $x^{\star}$ is a constrained Nash equilibrium of $\Gamma^{\epsilon}$.
@@
i.e., we have shown that $x^{\star}$ is a constrained Nash equilibrium of the original unperturbed game. ∎
@@ E.2 Technical lemma
Appendix F Proofs for Section 4
In this section, we provide the full proofs for the results of Section 4.
@@ F.1 Mixed-strategy potential games
In Example 3 we claimed that when considering mixed strategies over a finite action potential game, the potential structure is lifted to the continuous action game over compact convex set of mixed strategies. Here we state the corresponding proposition.
@@
Let $\mathcal{A}\coloneqq\times_{i\in[m]}\mathcal{A}_{i}$ with each $\mathcal{A}_{i}$ finite, and suppose $\widetilde{u}_{i}:\mathcal{A}\to[0,1]$ are player utilities and $\widetilde{c}_{j}:\mathcal{A}\to[0,1]$ constraints with thresholds $\alpha_{j}$ inducing feasible region $\widetilde{\mathcal{C}}$, such that there exists a potential function $\phi:\widetilde{\mathcal{C}}\to\mathbb{R}$ satisfying
$\phi(a)-\phi(a_{i}^{\prime},a_{-i})=\widetilde{u}_{i}(a)-\widetilde{u}_{i}(a_{i}^{\prime},a_{-i})\quad\text{for all }i\in[m],\,a\in\widetilde{\mathcal{C}},\,a_{i}^{\prime}\in\widetilde{\mathcal{C}}_{i}(a_{-i}).$
Then, the continuous action game with strategy space $\mathcal{X}\coloneqq\times_{i\in[m]}\Delta(\mathcal{A}_{i})$, utilities
@@
and constraints
$c_{j}(x)\coloneqq\mathbb{E}_{a\sim x}[\widetilde{c}_{j}(a)]$
with thresholds $\alpha_{j}$ is a constrained potential game with potential function $\Phi(x)\coloneqq\mathbb{E}_{a\sim x}[\phi(a)]$.
For the case of unconstrained potential games, Proposition 2 is a direct consequence of Monderer and Shapley, (1996), Lemma 2.10. The same reasoning extends straightforwardly to the constrained case; we therefore omit the proof here.
@@ F.2 Technical lemmas
If we have a game with utilities $\{u^{\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime}$, and another game with utilities $\{u^{\prime\prime}_{i}\}_{i\in[m]}$ and potential function $\Phi^{\prime\prime}$, it is immediate that then $\Phi^{\prime}+\Phi^{\prime\prime}$ is a potential function for the game with utilities $\{u^{\prime}_{i}+u^{\prime\prime}_{i}\}_{i\in[m]}$. The statement hence follows by observing that the sum of log barrier terms forms a cooperative game and hence $\eta\sum_{j\in[b]}\log(c_{j}(x)-\alpha_{j})$ is a potential function for it. ∎
@@ F.4 Lower bounding the stepsize
[…]